OpenStudy (anonymous):

1st order diff eq.: (dy)/(dx) - y/x = xe^x ; y(1) = e - 1

OpenStudy (lgbasallote):

cool linear DE

OpenStudy (lgbasallote):

@sami-21 your turn

OpenStudy (anonymous):

.....yaaaay..... jerkface... jk

OpenStudy (lgbasallote):

your in for a lot more integration :p >:)))

OpenStudy (anonymous):

mew!!!! or mu!

OpenStudy (lgbasallote):

this time you do 3 integration >:)))

OpenStudy (anonymous):

yup! looking forward to learning it

OpenStudy (lgbasallote):

....if you're lucky

OpenStudy (anonymous):

....noooo, i should be finding an integrating factor,.....

OpenStudy (lgbasallote):

sami's taking a long time to type something...

OpenStudy (lgbasallote):

must be a latex bomb

OpenStudy (anonymous):

.....life story?

OpenStudy (lgbasallote):

could be

OpenStudy (anonymous):

ok . first find the integrating factor. \[\Large e^{\int\limits_{}^{}\frac{-1}{x}dx} = e^{-\ln(x)}=e^{\ln(x)^{-1}}=x^{-1}=\frac{1}{x}\]

OpenStudy (anonymous):

multiply the whole equation with integrating factor.

OpenStudy (anonymous):

wait, how did you get the first part?

OpenStudy (anonymous):

are you familiar with linear equation ?

OpenStudy (anonymous):

if shown an example, yes

OpenStudy (anonymous):

i'm a visual guy

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

ok look this at site http://www.sosmath.com/diffeq/first/lineareq/lineareq.html

OpenStudy (anonymous):

P(x)y = y/x = (1/x)y :::: mew(x) = e^(integral: dx/x)

OpenStudy (anonymous):

yes it is -(1/x y) so e^integral (-1/x)

OpenStudy (anonymous):

forgot the negative, thank you

OpenStudy (anonymous):

ok, so now i multiply that to the whole equation?

OpenStudy (anonymous):

im, trying to do that math, but i am having no luck

OpenStudy (anonymous):

d/(dx)[e^(-Ln|x|)y] = xe^x * e^-ln|x| ???? whaaat?

OpenStudy (anonymous):

hmm its not that hard :P just multiply this and you will have 1/x(dy/dx)-y/x^2=e^x is it ok now ??

OpenStudy (anonymous):

now how did the right side come out to only e^x?

OpenStudy (anonymous):

x*1/x= ????????????????????

OpenStudy (anonymous):

and the step before that?

OpenStudy (anonymous):

which one ????

OpenStudy (anonymous):

why are you dividing it by x?

OpenStudy (anonymous):

i said multiply the whole equation with \[\frac{1}{x}\]

OpenStudy (anonymous):

ooooooh, i see now, e^-lnx = 1/x......

OpenStudy (anonymous):

ok, so how do i proceed from 1/x(dy/dx)-y/x^2=e^x ?

OpenStudy (anonymous):

put it all on the right side except for dy/dx and then solve for y?

OpenStudy (anonymous):

ok you can see the left the side is product rule of differentiation for d/dx(1/x*y)

OpenStudy (anonymous):

i did not see this...

OpenStudy (anonymous):

differentiate d/dx(1/x*y) use product rule. \[\Large \frac{d}{dx}(\frac{1}{x}y)=\frac{1}{x}\frac{dy}{dx}-\frac{1}{x^2}*y\] is it ok now ???????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the = e^x?

OpenStudy (anonymous):

i was just trying to show you the left side so it becomes \[\Large \frac{d}{dx}(\frac{1}{x}y)=e^x\] it it ok ??

OpenStudy (anonymous):

and then I integrate?

OpenStudy (anonymous):

or get y by itself again?

OpenStudy (anonymous):

yes integrate both sides . the left side derivative and integral will cancel each others. you will have \[\Large \frac{1}{x}*y=\int\limits_{}^{}e^x\] i hope you can do this from here :)))))))

OpenStudy (anonymous):

it wouldnt be easier to get y by itself?

OpenStudy (anonymous):

but would I get -2x^(-2) * (1/2)y^2 = e^x?

OpenStudy (anonymous):

or would i use the product rule again?

OpenStudy (anonymous):

since you got \[\Large \frac{1}{x}*y=\int\limits_{}^{}e^x\] just you need to integrate right side only \[\Large \frac{1}{x}y=e^x+C\] or by multiplying x \[\Large y=xe^x+Cx\]

OpenStudy (anonymous):

oye! so, after you multiply mew back in, how did (1/x)(dy/dx)−(y/x^2) = e^x become (dy/dx)((1/x)*y) = e^x

OpenStudy (anonymous):

oh, okay......

OpenStudy (anonymous):

because it was the result of product rule so i wrote (dy/dx)((1/x)*y) = e^x the left side as

OpenStudy (anonymous):

so then e-1 = e^x + c?

OpenStudy (anonymous):

you do not know integration of e^x ???

OpenStudy (anonymous):

i thought the integral of e^x WAS e^x...

OpenStudy (anonymous):

yes it is .

OpenStudy (anonymous):

i was plugging in the initial value

OpenStudy (anonymous):

ok plug in in the y=xe^x+Cx e^-1=e+C C=e^-1-e Or C=1/e-e Or C=(1-e^2)/e

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Thanks :P

OpenStudy (lgbasallote):

woah this is still ongoing? :o

OpenStudy (anonymous):

several breaks